OpenStudy (anonymous):

Find the value. 3/1 A. 1 B. 2 C. 3 D. 0.5

OpenStudy (anonymous):

@princed

OpenStudy (michele_laino):

what is 3:1=?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (danjs):

x * x * x = 1

OpenStudy (danjs):

\[\sqrt[n]{1} = 1~~~for ~all ~n\]

OpenStudy (danjs):

probably

OpenStudy (michele_laino):

what is the number which multiplied three times itself gives 1?

OpenStudy (anonymous):

Find the value

OpenStudy (michele_laino):

please what is: 1*1*1=?